Fermi National Laboratory


Accelerator Update for August 19, 2005 - August 22, 2005

Friday, August 19
The day shift began with Operations monitoring store 4341 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter and MTest experiments taking beam.

At 8:40 AM, a Tevatron (TeV) vacuum roughing pump (B4RP3) was found off. Per expert instructions, an operator turned it on, but the pump never got the vacuum good enough for the turbo pump to turn on. The pump valves were left closed.

At 9:56 AM, Operations turned off Booster to disconnect a RF station (BRF4) from the East Anode Power Supply (EAPS), but then the station turned on and ran normally, so nothing was done.

At 10:02 AM, Operations turned off the Main Injector (MI) so an expert could check out a Lambertson (LAM52). He swapped out a module in its power supply.

At 10:53 AM, Operations reported that two Linac RF stations went way out of phase. Linac experts adjusted them back to normal operation.

NuMI experts completed conducting target scans at 11:39 AM.

At 7:12 PM, operators reported problems with Pbar emittance blow up and dropping stacking rates. At 7:37 PM, operators halted stacking for a few minutes. Operators contacted several experts for help. At 8:25 PM, there was another emittance blow up.

Operations terminated store 4341 at 9:29 PM.

Operations put the TeV into shot setup at 9:49 PM.

Saturday, August 20
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

Operations established combination store 4342 at 12:01 AM with an initial luminosity of 100.71E30. Pbar resumed stacking.

Operations established beam to MCenter at 12:20 AM.

At 2:21 AM, a thunderstorm rolled over the lab and caused store 4342 to abort, then continuing lightning strikes caused the TeV to quench at sectors DA, DB, and DD. Booster and MI also suffered various system trips as the storm continued. Due the amount of trips in MI, the accelerator was turned off for about 30 minutes.

At 2:35 AM, Operations reported a network communication problem at sector C3. Operators tried unsuccessfully to reset the QPM. Operators called in an expert.

Pbar resumed stacking at 2:42 AM and NuMI resumed taking beam a minute later.

At 6:26 AM, Operations reported that there were still a few vacuum pumps off in Booster. Operators couldn't reset them remotely or locally. An expert will investigate this problem later.

Operations put the Recycler into shot setup to stash antiprotons at 7:50 AM.

Operations prepared CDF and D0 for access into their collision halls at 8:10 AM.

The Recycler stash ended and Pbar resumed stacking and MCenter resumed taking beam at 9:03 AM.

CDF ended their access at around 10:15 AM. D0 never went in on access.

Operations began the TeV checkout at 10:52 AM.

Operations put the Recycler into shot setup at 1:03 PM.

At 1:41 PM, Operations reported that the Recycler had lost its stash due to an MI interlock safety trip that caused its coasting beam valves to close. Operators determined that it was a PLC fault at MI52. Operations called in interlock system experts.

Operations put the TeV through a dry squeeze at 4:15 PM.

Operators reported at 5:14 PM that they had replaced three cables and two power supplies for Booster vacuum pumps and that they were all running normal.

At 5:46 PM, while interlocking MI, operators noticed a LCW leak from a quadrupole magnet in the P150 line. It was a slow leak so they didn't do anything except note it.

Interlock experts reported from MI at 7:57 PM that they had to change the PLC backplane along with its processor and power supply.

Operators had to re-interlock MI. At 9:02 PM, operators reported that the slow leak on the MI quadrupole had turned into a sprayer. An operator called in water experts.

Water experts completed a temporary repair on the quadrupole leak at 11:34 PM.

Pbar resumed stacking at 11:59 PM.

Sunday, August 21
The midnight shift began with the TeV in a dry squeeze, with Pbar stacking, and with MiniBooNE, NuMI taking beam.

MCenter resumed taking beam at 12:08 AM.

Operations reported at 12:45 AM that they had begun a proton only squeeze to tune the accelerator.

Operations loaded a 12x0 study-store for intrabeam scattering at 2:14 AM. Operators terminated the store two hours later.

Operations put the TeV into shot setup at 4:30 AM.

Operations established store 4345 at 6:41 AM with an initial luminosity of 57.1E30. Pbar resumed stacking and MCenter resumed taking beam.

Operations put the Recycler into shot setup at 11:15 AM.

The Recycler stash ended and Pbar resumed stacking and MCenter resumed taking beam at 12:18 PM.

Operations put the Recycler into shot setup at 4:26 PM.

The stash ended and Pbar and MCenter resumed normal operations at 5:23 PM.

Monday, August 22
The midnight shift began with Operations monitoring store 4345 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4345 at 4:07 AM.

Operations put the TeV into shot setup at 4:21 AM.

At 5:36 AM, the Booster EAPS tripped off due to an AC overload trip. Operators had it back up ten minutes later, but BRF4 was making such horrible sounds that they turned the station off. EAPS tripped off again at 5:58 AM. Operators reset it a couple of times and managed to get it up and running by 6:14 AM. BRF4 was then brought back on line.

Operations established combination store 4346 at 6:15 AM with an initial luminosity of 100.5E30. Pbar and MCenter resumed normal operations.

The Plans for Monday
The plans for today are to store, stack, and stash.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/22/2005   email Fermilab